OpenStudy (vera_ewing):

Is anyone good at Algebra 2?

OpenStudy (vera_ewing):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = –1.

OpenStudy (vera_ewing):

@whpalmer4 I'm pretty sure the answer is f(x) = 4x^2 right?

OpenStudy (anonymous):

equidistant meaning from the vertex to the focus, is the same distance as from the vertex to the directrix what does that mean? well, the vertex is half-way between both of those

OpenStudy (vera_ewing):

Ok so I'm right? My answer is correct?

OpenStudy (anonymous):

half-way in between those 2 points... .if we keep on going up from the focus, notice that we're just moving between (2,−1)→(2,−12) ↑half-way in between Yes , you are correct! :)

OpenStudy (vera_ewing):

Yes! Thank you! :) @samanthha

OpenStudy (xapproachesinfinity):

I'm not convinced hehe

OpenStudy (anonymous):

im sorry about that, it didn't turn out like it was supposed to, but you are correct!

OpenStudy (vera_ewing):

@xapproachesinfinity WHAT?!??!

OpenStudy (vera_ewing):

@xapproachesinfinity So I'm wrong?!?

OpenStudy (xapproachesinfinity):

i said that was not a convincing answer to me! that's all!! you can show all the work to why the answer must be that?

OpenStudy (vera_ewing):

@xapproachesinfinity You must know the answer to that question? ;) Great! Now you've made me feel guilty :'(

OpenStudy (xapproachesinfinity):

eh why! i'm just not sure you answered right! if you feel it is right then that's okay with me

OpenStudy (vera_ewing):

@xapproachesinfinity So there is a possiblity that I'm wrong?!!!

OpenStudy (xapproachesinfinity):

hmm i would do this first \(\huge f(x)=a(x-h)^2+k\) where (h,k) is the vertex find the vertex by midpoint forumla

OpenStudy (xapproachesinfinity):

according to you if the equation is \(\huge f(x)=4x^2=4(x-0)^2+0\) so the vertex is at (0,0) is that true?

OpenStudy (vera_ewing):

@xapproachesinfinity omg

OpenStudy (xapproachesinfinity):

hhehe actually you are correct!

OpenStudy (xapproachesinfinity):

but you didn't know how to justify it

OpenStudy (anonymous):

I'm sorry, i should of actually explained.

OpenStudy (vera_ewing):

@xapproachesinfinity I can honestly say I'm a great guesser lol

OpenStudy (xapproachesinfinity):

hehehe no guesses okay

OpenStudy (anonymous):

actually it is wrong

OpenStudy (vera_ewing):

@jkasper are you kidding me???

OpenStudy (anonymous):

it should be y=(1/4) x^2

OpenStudy (vera_ewing):

@jkasper are you sure?